# Practical hybrid decoding scheme for parity-encoded spin systems

### v3

@@ II.2 Parity encoding architecture and LDPC codes
We illustrate a PE architecture with concrete examples. The PE architecture corresponds to the classical ECC according to the following map: $\bar{z}_{ij}=\bar{Z}_{i}\oplus\bar{Z}_{j}$ for $1\leq i\leq j\leq k$. As a simple example, consider the case $K=4$ and assume that $\boldsymbol{\bar{Z}}=\left(\bar{Z}_{1},\bar{Z}_{2},\bar{Z}_{3},\bar{Z}_{4}\right)\in\left\{0,1\right\}^{K}$ and $\bar{\boldsymbol{z}}=\left(\bar{z}_{12},\bar{z}_{13},\bar{z}_{14},\bar{z}_{23},\bar{z}_{24},\bar{z}_{34}\right)\in\left\{0,1\right\}^{\tbinom{K}{2}}$. Because each element of $\bar{\boldsymbol{z}}$ is the binary sum of two elements in $\boldsymbol{\bar{Z}}$, the generating matrix associated with this map is given by the $K\times\tbinom{K}{2}$ matrix:
$\boldsymbol{G}=\begin{pmatrix}1&1&1&0&0&0\\ 1&0&0&1&1&0\\ 0&1&0&1&0&1\\ 0&0&1&0&1&1\end{pmatrix}$. (1)
@@
Note that there are two 1s in each column in $\boldsymbol{G}$. Then, we can choose the following two parity check matrices,
$\boldsymbol{H}=\begin{pmatrix}1&1&0&1&0&0\\ 0&1&1&1&1&0\\ 0&0&0&1&1&1\end{pmatrix}$ (2)
@@
$\boldsymbol{H}^{\prime}=\begin{pmatrix}1&1&0&1&0&0\\ 1&0&1&0&1&0\\ 0&1&1&0&0&1\\ 0&0&0&1&1&1\end{pmatrix}$, (3)
@@
satisfying the constraints $\boldsymbol{G}\boldsymbol{H}^{T}=\boldsymbol{G}\boldsymbol{H}^{\prime}{}^{T}=\boldsymbol{0}$. From these matrices, two different syndrome vectors can be derived. One is weight-4 syndrome vector $\boldsymbol{\bar{s}}^{(4)}(\bar{\boldsymbol{x}})=\left(\bar{s}_{1223}^{(4)},\bar{s}_{1234}^{(4)},\bar{s}_{2334}^{(4)}\right)=\bar{\boldsymbol{x}}\boldsymbol{H}^{T}\in\left\{0,1\right\}^{\tbinom{K-1}{2}}$, and the other is weight-3 syndrome vector $\boldsymbol{\bar{s}}^{(3)}(\bar{\boldsymbol{x}})=\left(\bar{s}_{123}^{(3)},\bar{s}_{124}^{(3)},\bar{s}_{134}^{(3)},\bar{s}_{234}^{(3)}\right)=\bar{\boldsymbol{x}}\boldsymbol{H}^{\prime}{}^{T}\in\left\{0,1\right\}^{\tbinom{K}{3}}$, where $\bar{\boldsymbol{x}}=\left(\bar{x}_{12},\bar{x}_{13},\bar{x}_{14},\bar{x}_{23},\bar{x}_{24},\bar{x}_{34}\right)\in\left\{0,1\right\}^{\tbinom{K}{2}}$ is an arbitrary binary vector. $\boldsymbol{\bar{s}}^{(3)}(\bar{\boldsymbol{x}})$ is weight-3 because each of its elements is written as a linear combination of three elements in $\bar{\boldsymbol{x}}$: $\bar{s}_{klm}^{(3)}=\bar{x}_{kl}\oplus\bar{x}_{lm}\oplus\bar{x}_{km}$, where $1\leq k<l<m\leq K$. Similarly, $\boldsymbol{\bar{s}}^{(4)}(\bar{\boldsymbol{x}})$ is weight-4 because each of its elements is written as a linear combination of four elements in $\bar{\boldsymbol{x}}$ if $\bar{\boldsymbol{x}}$ is complemented by the fictitious components $\bar{x}_{ii}\,(i=2,\ldots,K-1)$ assigned by a fixed value 0: $\bar{s}_{klmn}^{(4)}=\bar{x}_{kl}\oplus\bar{x}_{lm}\oplus\bar{x}_{ln}\oplus\bar{x}_{kn}$. We note that any element in $\boldsymbol{\bar{s}}^{(4)}(\bar{\boldsymbol{x}})$ can be written as a linear combination of appropriate elements in $\boldsymbol{\bar{s}}^{(3)}(\bar{\boldsymbol{x}})$, and vice versa.
@@
The connections between the variables $\bar{x}_{kl}$ and the weight-4 syndromes $\bar{s}_{klmn}^{(4)}$ are depicted by a sparse bipartite graph shown in Fig.2(a). Similar connections between variables $\bar{x}_{kl}$ and the weight-3 syndromes $\bar{s}_{klm}^{(3)}$ are depicted in Fig.2(b). The number of elements in $\bar{\boldsymbol{x}}$ is $N_{v}=\tbinom{K}{2}$, while the number of elements in $\boldsymbol{\bar{s}}^{(4)}(\bar{\boldsymbol{x}})$ and $\boldsymbol{\bar{s}}^{(3)}(\bar{\boldsymbol{x}})$ is $N_{c}=\tbinom{K-1}{2}$ and $N_{c}=\tbinom{K}{3}$, respectively. In the terminology of graph theory, the $N_{v}$ elements in $\bar{\boldsymbol{x}}$ constitute variable nodes (VNs) and the $N_{c}$ elements in $\boldsymbol{\bar{s}}^{(4)}(\bar{\boldsymbol{x}})$ or $\boldsymbol{\bar{s}}^{(3)}(\bar{\boldsymbol{x}})$ constitute check nodes (CNs). The matrix $\boldsymbol{H}$ or $\boldsymbol{H}^{\prime}$ has $N_{c}$ rows and $N_{v}$ columns, where each row $i$ represents a CN and each column $j$ represents a VN; If the entry $H_{ij}=1$, VN $j$ is connected to CN $i$ by an edge. Thus, each edge connecting VN and CN corresponds to an entry $1$ in the row and column of $\boldsymbol{H}$ or $\boldsymbol{H}^{\prime}$. Let $d_{c}$ be the number of $1$s in each row and $d_{v}$ be the number of $1$s in each column of $\boldsymbol{H}$ or $\boldsymbol{H}^{\prime}$. These numbers represent the number of edges connected to a VN and a CN, respectively, referred to as row and column weights. The matrix $\boldsymbol{H}^{\prime}$ is regular, where the row weight $d_{c}=3$ is common for all the CNs and the column weight $d_{v}=K-2$ is common for all the VNs. In contrast, the matrix $\boldsymbol{H}$ is irregular, and its weights depend on the associated nodes, which are at most 4.
@@
If $K$ is very large, we can always choose a sparse matrix $\boldsymbol{H}$ such that most of the entries are 0 and a few ones are $1$. Then, the PE architecture is closely connected to the LDPC codes. Only $2^{K}$ elements out of the $2^{\tbinom{K}{2}}$ possible $\bar{\boldsymbol{x}}$ are valid code-words and satisfy the parity check constraints $\boldsymbol{\bar{s}}^{(4)}(\bar{\boldsymbol{x}})=\boldsymbol{0}_{1\times\tbinom{K-1}{2}}$ or $\boldsymbol{\bar{s}}^{(3)}(\bar{\boldsymbol{x}})=\boldsymbol{0}_{1\times\tbinom{K}{3}}$. Conversely, the syndrome vector $\boldsymbol{\bar{s}}^{(n)}(\bar{\boldsymbol{x}})$ should be all-zero vector, i.e., the vector whose elements are all zero, for the associated vector $\bar{\boldsymbol{x}}$ to be a valid code-word.
@@
Table 1: Parameters related to Fig.2 for $4\leq K\leq 7$.
@@ II.3 SLHZ model
In the remainder of this paper, we will discuss our arguments primarily in the language of spin glass. Following Sourlas, our argument relies on isomorphism between the additive Boolean group $\left(\left\{0,1\right\},\oplus\right)$ and the multiplicative Ising group $\left(\left\{\pm 1\right\},\cdot\right)$, where a binary variable $\bar{a}{}_{i}\in\left\{0,1\right\}$ maps to the spin variable $a_{i}=\left(-1\right)^{\bar{a}_{i}}\in\left\{\pm 1\right\}$ and the binary sum maps to the product by $a_{i}a_{j}=(-1)^{\bar{a}_{i}\oplus\bar{a}_{j}}\in\left\{\pm 1\right\}$ [5, 6, 7, 25]. Source-word $\bar{\boldsymbol{Z}}$ and code-word $\bar{\boldsymbol{z}}$ are mapped to vectors $\boldsymbol{Z}=\left(Z_{1},\ldots,Z_{K}\right)\in\left\{\pm 1\right\}^{K}$ and $\boldsymbol{z}=\left(z_{1},\ldots,z_{N_{v}}\right)\in\left\{\pm 1\right\}^{N_{v}}$ in the spin representation, respectively. Hereafter, we will refer to $\boldsymbol{Z}$ and $\boldsymbol{z}$ as a source-state and a code-state associated with source-word $\bar{\boldsymbol{Z}}$ and code-word $\bar{\boldsymbol{z}}$, respectively, and associated fictitious spins as logical and physical spins, respectively. Similarly, binary vectors $\bar{\boldsymbol{s}}$ and $\bar{\boldsymbol{x}}$ are mapped to the associated vectors $\boldsymbol{s}$ and $\boldsymbol{x}$ in the spin representation. In the following, we denote variables by symbols with and without an overbar in the binary and spin representations. It should be noted that by isomorphism, every addition of two binary variables corresponds to a unique product of spin variables and vice versa. For example, since $\bar{\boldsymbol{z}}=\boldsymbol{\bar{Z}}\boldsymbol{G}\>\left(\mathrm{mod}\>2\right)$ holds, $\boldsymbol{z}$ and $\boldsymbol{Z}$ are connected by the relation
@@
for $i=1,\ldots,N_{c}$, if $\bar{\boldsymbol{z}}$ satisfies the parity check equation $\bar{\boldsymbol{z}}\boldsymbol{H}^{T}\>\left(\mathrm{mod}\>2\right)=\boldsymbol{0}_{1\times N_{c}}$.
Now, let us consider the graph in Fig.3, which is topologically equivalent to Fig.2(a). This graph is nothing but the SLHZ model [10]. This implies that the SLHZ model is essentially a PE-based model that utilizes the weight-4 syndrome, as viewed from the perspective of the spin glass model. The plaquette referred to by LHZ [10] and the stabilizer by A. Rocchetto, S. C. Benjamin, and Y. Li [26] corresponds to the weight-4 syndrome from the perspective of the classical ECCs, each of the elements is given by a product of four elements in $\boldsymbol{x}$ if $\boldsymbol{x}$ is complemented by the fictitious spin variables $x_{ii}\,(i=2,\ldots,K-1)$ with fixed value +1+1: $s_{klmn}^{(4)}(\boldsymbol{x})=x_{km}x_{lm}x_{ln}x_{kn}$ [10]. The vector $\boldsymbol{s}^{(n)}(\boldsymbol{x})$ should be an all-one vector, i.e., the vector whose elements are all one, for the associated state $\boldsymbol{x}$ of the physical spin to be a valid code-state.
Figure 3: Bipartite graph topologically equivalent to Fig.2(a). This graph avoids any edge crossings.
III Post-readout decoding of the SLHZ model
In this section, we consider post-readout decoding of the SLHZ model from the perspective of the classical ECCs. We first present a general probabilistic framework for classical decoding of ECCs. Next, we propose a simple, straightforward decoding algorithm tailored for the SLHZ model.
@@ III.1 Probabilistic decoding
Post-readout decoding can be performed according to Bayesian inference. Consider the conditional probability $P\left(\boldsymbol{z}|\boldsymbol{y}\right)d\boldsymbol{y}$ that the prepared code-state is $\boldsymbol{z}$ when the observation was between $\boldsymbol{y}$ and $\boldsymbol{y}+d\boldsymbol{y}$. According to Bayes’ theorem,
$P(\boldsymbol{z}|\boldsymbol{y})=\frac{P(\boldsymbol{y}|\boldsymbol{z})P(\boldsymbol{z})}{\sum_{\boldsymbol{z}}P(\boldsymbol{y}|\boldsymbol{z})P(\boldsymbol{z})}=\kappa P(\boldsymbol{y}|\boldsymbol{z})P(\boldsymbol{z})$ (6)
holds, where $\kappa$ is a constant independent of $\boldsymbol{z}$ which is determined by the normalization condition $\mathop{\sum_{\boldsymbol{z}}P\left(\boldsymbol{z}|\boldsymbol{y}\right)=1}$, and $P\left(\boldsymbol{z}\right)$ is the prior probability for the code-state $\boldsymbol{z}$. According to Bayesian inference, we can consider the state that maximizes the conditional probability $P\left(\boldsymbol{z}|\boldsymbol{y}\right)$, which offers the most probable word (“word maximum a posteriori probability” or “word MAP decoding”). Alternatively, we can consider the state that maximizes its marginals, which offers the most probable symbol (“symbol maximum a posteriori probability” or “symbol MAP decoding”).
Word MAP decoding
We follow the argument given by Sourlas, who first pointed out the link between the ECCs and the spin glass model [4]. Sourlas showed that, based on Eqs.(4) and (5), two different formulations are possible for word MAP decoding of the same ECC. The first formulation is given in terms of source-state $\boldsymbol{Z}$. In this formulation, we assume the following prior probability for $\boldsymbol{z}$:
$P(\boldsymbol{z})=\mu\prod_{i=1}^{N_{v}}\delta\left(z_{i},\prod_{\{j:G_{ji}=1\}}Z_{j}\right),$ (7)
where $\mu$ is a normalization constant. The Kronecker’s $\delta$’s in Eq.(7) enforce the constraint that $\boldsymbol{z}$ obeys Eq.(4); it is a valid code-state. Assuming that the noise is independent for each spin and that $P\left(\boldsymbol{y}|\boldsymbol{z}\right)=\mathop{\underset{i=1}{\stackrel{{\scriptstyle N_{v}}}{{\prod}}}P\left(y_{i}|z_{i}\right)}$ holds (memoryless channel), we can derive the following equation:
$-\ln P(\boldsymbol{z}|\boldsymbol{y})$ $=$ $\mathrm{const.}-\sum_{i=1}^{N_{v}}B_{i}\prod_{\{j:G_{ji}=1\}}Z_{j}$ (8)
$\equiv$ $H^{source}(\boldsymbol{Z}),$
@@
is the half log-likelihood ratio (LLR) for the channel observation $y_{i}$ [5]. The vector $\boldsymbol{B}=\left(B_{1},\ldots,B_{N_{v}}\right)\in\mathbb{R}^{N_{v}}$ contains all the information about the observation $\boldsymbol{y}$. In this paper, we focus on the communication through the additive white Gaussian noise (AWGN) channel. The likelihood is given as
@@
where $\left|v\right|$ and $\sigma^{2}$ are the amplitude of the prepared signal and the variance of the common Gaussian noise. Then the LLR is given by $B_{i}=\tfrac{1}{2}\beta y_{i}$, where $\beta=\tfrac{2|v|}{\sigma^{2}}>0$ is called the channel reliability factor, and its inverse $\beta^{-1}$ is corresponds to the temperature in the language of spin glasses. Note that $\frac{1}{2}$$\left(\frac{|v|}{\sigma}\right)^{2}$ is the signal-to-noise ratio (SNR) of the AWGN channel. Thus, $B_{i}$ is proportional to the magnitude of the channel observation $y_{i}$ for the AWGN channel. It is evident that $H^{source}(\boldsymbol{Z})$ is in the form of a spin glass Hamiltonian, where $\tfrac{1}{2}y_{i}$ is identified with the coupling constant $J_{i}$. In this formulation, the word MAP decoding corresponds to finding the ground state of the Hamiltonian $H^{source}(\boldsymbol{Z})$.
Alternatively, the second formulation is given in terms of code-state $\boldsymbol{z}$. In this formulation, according to Eq.(5), we assume the following prior probability for $\boldsymbol{z}$:
$P(\boldsymbol{z})=\mu\prod_{i=1}^{N_{c}}\delta(s_{i}\left(\boldsymbol{z}\right),+1),$ (11)
@@
$s_{i}(\boldsymbol{z})=\prod_{\{j:H_{ij}=1\}}z_{j}$ (12)
is the $i$th syndrome for $\boldsymbol{z}$ in the spin representation. Then, we can derive the following equation:
$-\ln P(\boldsymbol{z}|\boldsymbol{y})$ $=$ $-\sum_{i=1}^{N_{v}}B_{i}z_{i}+\underset{\gamma\rightarrow\infty}{\lim}\gamma\sum_{i=1}^{N_{c}}\frac{1-s_{i}(\boldsymbol{z})}{2}$ (13)
$\equiv$ $H^{code}(\boldsymbol{z}).$
In Eq.(13), $\delta$’s in Eq.(11) is replaced by a soft constraint using the identity

### v4

@@ II.2 Parity-encoding architecture and LDPC codes
We illustrate a PE architecture with concrete examples. The PE architecture corresponds to the classical ECC according to the following map: $\bar{z}_{ij}=\bar{Z}_{i}\oplus\bar{Z}_{j}$ for $1\leq i\leq j\leq k$. As a simple example, consider the case $K=4$ and assume that $\bar{\textbf{Z}}=\left(\bar{Z}_{1},\bar{Z}_{2},\bar{Z}_{3},\bar{Z}_{4}\right)\in\left\{0,1\right\}^{K}$ and $\bar{\textbf{z}}=\left(\bar{z}_{12},\bar{z}_{13},\bar{z}_{14},\bar{z}_{23},\bar{z}_{24},\bar{z}_{34}\right)\in\left\{0,1\right\}^{\tbinom{K}{2}}$. Because each element of $\bar{\textbf{z}}$ is the binary sum of two elements in $\bar{\textbf{Z}}$, the generating matrix associated with this map is given by the $K\times\tbinom{K}{2}$ matrix:
$\textbf{G}=\begin{pmatrix}1&1&1&0&0&0\\ 1&0&0&1&1&0\\ 0&1&0&1&0&1\\ 0&0&1&0&1&1\end{pmatrix}$. (1)
@@
Note that there are two instances of 1 in each column in G. Then, we can choose the following two parity-check matrices,
$\textbf{H}=\begin{pmatrix}1&1&0&1&0&0\\ 0&1&1&1&1&0\\ 0&0&0&1&1&1\end{pmatrix}$ (2)
@@
$\textbf{H}^{\prime}=\begin{pmatrix}1&1&0&1&0&0\\ 1&0&1&0&1&0\\ 0&1&1&0&0&1\\ 0&0&0&1&1&1\end{pmatrix}$, (3)
@@
satisfying the constraints $\textbf{G}\textbf{H}^{T}=\textbf{G}\textbf{H}^{\prime}{}^{T}=\boldsymbol{0}$. From these matrices, two different syndrome vectors can be derived. One is weight-4 syndrome vector $\bar{\textbf{s}}^{(4)}(\bar{\textbf{x}})=\left(\bar{s}_{1223}^{(4)},\bar{s}_{1234}^{(4)},\bar{s}_{2334}^{(4)}\right)=\bar{\textbf{x}}\textbf{H}^{T}\in\left\{0,1\right\}^{\tbinom{K-1}{2}}$, and the other is weight-3 syndrome vector $\bar{\textbf{s}}^{(3)}(\bar{\boldsymbol{x}})=\left(\bar{s}_{123}^{(3)},\bar{s}_{124}^{(3)},\bar{s}_{134}^{(3)},\bar{s}_{234}^{(3)}\right)=\bar{\textbf{x}}\textbf{H}^{\prime}{}^{T}\in\left\{0,1\right\}^{\tbinom{K}{3}}$, where $\bar{\textbf{x}}=\left(\bar{x}_{12},\bar{x}_{13},\bar{x}_{14},\bar{x}_{23},\bar{x}_{24},\bar{x}_{34}\right)\in\left\{0,1\right\}^{\tbinom{K}{2}}$ is an arbitrary binary vector. $\bar{\textbf{s}}^{(3)}(\bar{\textbf{x}})$ is weight-3 because each of its elements is written as a linear combination of three elements in $\bar{\textbf{x}}$: $\bar{s}_{klm}^{(3)}=\bar{x}_{kl}\oplus\bar{x}_{lm}\oplus\bar{x}_{km}$, where $1\leq k<l<m\leq K$. Similarly, $\bar{\textbf{s}}^{(4)}(\bar{\textbf{x}})$ is weight-4 because each of its elements is written as a linear combination of four elements in $\bar{\textbf{x}}$ if $\bar{\textbf{x}}$ is complemented by the fictitious components $\bar{x}_{ii}\,(i=2,\ldots,K-1)$ assigned by a fixed value 0: $\bar{s}_{klmn}^{(4)}=\bar{x}_{kl}\oplus\bar{x}_{lm}\oplus\bar{x}_{ln}\oplus\bar{x}_{kn}$. We note that any element in $\bar{\textbf{s}}^{(4)}(\bar{\textbf{x}})$ can be written as a linear combination of appropriate elements in $\bar{\textbf{s}}^{(3)}(\bar{\textbf{x}})$, and vice versa.
@@
The connections between the variables $\bar{x}_{kl}$ and the weight-4 syndromes $\bar{s}_{klmn}^{(4)}$ are depicted by a sparse bipartite graph shown in Fig.2(a). Similar connections between variables $\bar{x}_{kl}$ and the weight-3 syndromes $\bar{s}_{klm}^{(3)}$ are depicted in Fig.2(b). The number of elements in $\bar{\textbf{x}}$ is $N_{v}=\tbinom{K}{2}$, while the numbers of elements in $\bar{\textbf{s}}^{(4)}(\bar{\textbf{x}})$ and $\bar{\textbf{s}}^{(3)}(\bar{\textbf{x}})$ are $N_{c}=\tbinom{K-1}{2}$ and $N_{c}=\tbinom{K}{3}$, respectively. In the terminology of graph theory, the $N_{v}$ elements in $\bar{\textbf{x}}$ constitute variable nodes (VNs) and the $N_{c}$ elements in $\bar{\textbf{s}}^{(4)}(\bar{\textbf{x}})$ or $\bar{\textbf{s}}^{(3)}(\bar{\textbf{x}})$ constitute check nodes (CNs). The matrix H or $\textbf{H}^{\prime}$ has $N_{c}$ rows and $N_{v}$ columns, where each row $i$ represents a CN and each column $j$ represents a VN; If the entry $H_{ij}=1$, VN $j$ is connected to CN $i$ by an edge. Thus, each edge connecting VN and CN corresponds to an entry $1$ in the row and column of H or $\textbf{H}^{\prime}$. Let $d_{c}$ be the number of $1$s in each row and $d_{v}$ be the number of $1$s in each column of H or $\textbf{H}^{\prime}$. These numbers represent the number of edges connected to a VN and a CN, respectively, referred to as row and column weights. The matrix $\textbf{H}^{\prime}$ is regular, where the row weight $d_{c}=3$ is common for all the CNs and the column weight $d_{v}=K-2$ is common for all the VNs. In contrast, the matrix H is irregular, and its weights depend on the associated nodes, which are at most 4.
@@
If $K$ is very large, we can always choose a sparse matrix H such that most of the entries are 0 and a few ones take value $1$. Then, the PE architecture is closely connected to the LDPC codes. Only $2^{K}$ elements out of the $2^{\tbinom{K}{2}}$ possible $\bar{\textbf{x}}$ are valid code words and satisfy the parity-check constraints $\bar{\textbf{s}}^{(4)}(\bar{\textbf{x}})=\boldsymbol{0}_{1\times\tbinom{K-1}{2}}$ or $\bar{\textbf{s}}^{(3)}(\bar{\textbf{x}})=\boldsymbol{0}_{1\times\tbinom{K}{3}}$. Conversely, the syndrome vector $\bar{\textbf{s}}^{(n)}(\bar{\textbf{x}})$ should be the all-zero vector, i.e., the vector the elements of which are all zero, for the associated vector $\bar{\textbf{x}}$ to be a valid code word.
@@
Table 1: The parameters related to Fig.2 for $4\leq K\leq 7$.
@@ II.3 SLHZ model
In the remainder of this paper, we will discuss our arguments primarily in the language of spin glasses. Following Sourlas, our argument relies on isomorphism between the additive Boolean group $\left(\left\{0,1\right\},\oplus\right)$ and the multiplicative Ising group $\left(\left\{\pm 1\right\},\cdot\right)$, where a binary variable $\bar{a}{}_{i}\in\left\{0,1\right\}$ maps to the spin variable $a_{i}=\left(-1\right)^{\bar{a}_{i}}\in\left\{\pm 1\right\}$ and the binary sum maps to the product by $a_{i}a_{j}=(-1)^{\bar{a}_{i}\oplus\bar{a}_{j}}\in\left\{\pm 1\right\}$ [5, 6, 7]. Source word $\bar{\textbf{Z}}$ and code word $\bar{\textbf{z}}$ are mapped to vectors $\textbf{Z}=\left(Z_{1},\ldots,Z_{K}\right)\in\left\{\pm 1\right\}^{K}$ and $\textbf{z}=\left(z_{1},\ldots,z_{N_{v}}\right)\in\left\{\pm 1\right\}^{N_{v}}$ in the spin representation, respectively. Hereafter, we will refer to Z and z as a source-state and a code-state associated with source word $\bar{\textbf{Z}}$ and code word $\bar{\textbf{z}}$, respectively, and associated fictitious spins as logical and physical spins, respectively. Similarly, binary vectors $\bar{\textbf{s}}$ and $\bar{\textbf{x}}$ are mapped to the associated vectors s and x in the spin representation. In the following, we denote variables by symbols with and without an overbar in the binary and spin representations. It should be noted that by isomorphism, every addition of two binary variables corresponds to a unique product of spin variables and vice versa. For example, since $\bar{\textbf{z}}=\bar{\textbf{Z}}\textbf{G}\>\left(\mathrm{mod}\>2\right)$ holds, z and Z are connected by the relation
@@
for $i=1,\ldots,N_{c}$, if $\bar{\textbf{z}}$ satisfies the parity-check equation $\bar{\textbf{z}}\textbf{H}^{T}\>\left(\mathrm{mod}\>2\right)=\boldsymbol{0}_{1\times N_{c}}$.
Now, let us consider the graph in Fig.3, which is topologically equivalent to Fig.2(a). This graph is simply the SLHZ model [10]. This implies that the SLHZ model is essentially a PE-based model that utilizes the weight-4 syndrome, as viewed from the perspective of the spin-glass model. The plaquette referred to by LHZ [10] and the stabilizer by Rocchetto, Benjamin, and Li [25] correspond to the weight-4 syndrome from the perspective of the classical ECCs. Each of the elements is given by a product of four elements in x if x is complemented by the fictitious spin variables $x_{ii}\,(i=2,\ldots,K-1)$ with fixed value +1+1: $s_{klmn}^{(4)}(\textbf{x})=x_{km}x_{lm}x_{ln}x_{kn}$ [10]. The vector $\textbf{s}^{(n)}(\textbf{x})$ should be an all-one vector, i.e., the vector the elements of which are all one, for the associated state x of the physical spin to be a valid code state.
Figure 3: The bipartite graph that is topologically equivalent to Fig.2(a). This graph avoids any edge crossings.
III POSTREADOUT DECODING OF THE SLHZ MODEL
In this section, we consider postreadout decoding of the SLHZ model from the perspective of the classical ECCs. We first present a general probabilistic framework for classical decoding of ECCs. Next, we propose a simple straightforward decoding algorithm tailored for the SLHZ model.
@@ III.1 Probabilistic decoding
Postreadout decoding can be performed according to Bayesian inference. Consider the conditional probability $P\left(\textbf{z}|\textbf{y}\right)d\textbf{y}$ that the prepared code state is z when the observation was between y and $\textbf{y}+d\textbf{y}$. According to the Bayes theorem,
$P(\textbf{z}|\textbf{y})=\frac{P(\textbf{y}|\textbf{z})P(\textbf{z})}{\sum_{\textbf{z}}P(\textbf{y}|\textbf{z})P(\textbf{z})}=\kappa P(\textbf{y}|\textbf{z})P(\textbf{z})$ (6)
holds, where $\kappa$ is a constant independent of z which is determined by the normalization condition $\mathop{\sum_{\textbf{z}}P\left(\textbf{z}|\textbf{y}\right)=1}$, and $P\left(\textbf{z}\right)$ is the prior probability for the code state z. According to Bayesian inference, we can consider the state that maximizes the conditional probability $P\left(\textbf{z}|\textbf{y}\right)$, which offers the most probable word (“word maximum a posteriori probability” or “word-MAP" decoding). Alternatively, we can consider the state that maximizes its marginals, which offers the most probable symbol (“symbol maximum a posteriori probability” or “symbol-MAP" decoding).
Word-MAP decoding
We follow the argument given by Sourlas, who first pointed out the link between the ECCs and the spin glass-model [4]. Sourlas showed that, based on Eqs.(4) and (5), two different formulations are possible for word-MAP decoding of the same ECC. The first formulation is given in terms of source state Z. In this formulation, we assume the following prior probability for z:
$P(\textbf{z})=\mu\prod_{i=1}^{N_{v}}\delta\left(z_{i},\prod_{\{j:G_{ji}=1\}}Z_{j}\right),$ (7)
where $\mu$ is a normalization constant. The Kronecker’s $\delta$’s in Eq.(7) enforce the constraint that $\boldsymbol{z}$ obeys Eq.(4); it is a valid code state. Assuming that the noise is independent for each spin and that $P\left(\textbf{y}|\textbf{z}\right)=\mathop{\underset{i=1}{\stackrel{{\scriptstyle N_{v}}}{{\prod}}}P\left(y_{i}|z_{i}\right)}$ holds (memoryless channel), we can derive the following equation:
$-\ln P(\boldsymbol{z}|\boldsymbol{y})$ $=$ $\mathrm{constant}-\sum_{i=1}^{N_{v}}B_{i}\prod_{\{j:G_{ji}=1\}}Z_{j}$ (8)
$\equiv$ $H^{source}(\textbf{Z}),$
@@
is the half log-likelihood ratio (LLR) for the channel observation $y_{i}$ [5]. The vector $\textbf{B}=\left(B_{1},\ldots,B_{N_{v}}\right)\in\mathbb{R}^{N_{v}}$ contains all the information about the observation $\boldsymbol{y}$. In this paper, we focus on the communication through the AWGN channel. The likelihood is given as
@@
where $\left|v\right|$ and $\sigma^{2}$ are the amplitude of the prepared signal and the variance of the common Gaussian noise. Then the LLR is given by $B_{i}=\tfrac{1}{2}\beta y_{i}$, where $\beta=\tfrac{2|v|}{\sigma^{2}}>0$ is called the channel reliability factor, and its inverse $\beta^{-1}$ corresponds to the temperature in the language of spin glasses. Note that $\frac{1}{2}$$\left(\frac{|v|}{\sigma}\right)^{2}$ is the signal-to-noise ratio (SNR) of the AWGN channel. Thus, $B_{i}$ is proportional to the magnitude of the channel observation $y_{i}$ for the AWGN channel. It is evident that $H^{source}(\textbf{Z})$ is in the form of a spin-glass Hamiltonian, where $\tfrac{1}{2}y_{i}$ is identified with the coupling constant $J_{i}$. In this formulation, the word-MAP decoding corresponds to finding the ground state of the Hamiltonian $H^{source}(\textbf{Z})$.
Alternatively, the second formulation is given in terms of code-state z. In this formulation, according to Eq.(5), we assume the following prior probability for z:
$P(\textbf{z})=\mu\prod_{i=1}^{N_{c}}\delta(s_{i}\left(\textbf{z}\right),+1),$ (11)
@@
$s_{i}(\textbf{z})=\prod_{\{j:H_{ij}=1\}}z_{j}$ (12)
is the $i$th syndrome for z in the spin representation. Then, we can derive the following equation:
$-\ln P(\textbf{z}|\textbf{y})$ $=$ $-\sum_{i=1}^{N_{v}}B_{i}z_{i}+\underset{\gamma\rightarrow\infty}{\lim}\gamma\sum_{i=1}^{N_{c}}\frac{1-s_{i}(\textbf{z})}{2}$ (13)
$\equiv$ $H^{code}(\textbf{z}).$
In Eq.(13), the $\delta$’s in Eq.(11) are replaced by a soft constraint using the identity